OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x^2 = −6x − 7?

OpenStudy (anonymous):

y=x^2-6x-7 x^2-6x-7=0 (x-7)(x+1)=0 x=7 or x=-1

OpenStudy (anonymous):

(1, 0) and (7, 0) (1, −7) (−1, 0) and (7, 0) (−1, 0) and (−7, 0)

OpenStudy (anonymous):

answer is third one. do you understand?

OpenStudy (anonymous):

yeah thnx

OpenStudy (anonymous):

welcome :)